How many grams of magnesium oxide are obtained during the combustion of 1 g of magnesium in air according to the reaction 2Mg (s) + O2 (g) -> 2MgO (s). Ar(Mg) = 24.3; Ar(O) = 16



1
Expert's answer
2021-09-20T02:26:59-0400

Moles of "Mg=\frac{2}{24.3}=0.082moles"


Ratio of "Mg" and "MgO= 2:2"


Hence moles of of "MgO=" "0.082moles"


Molar mass of "MgO=24.3+16=40.3"


Mass of "MgO=40.3\u00d70.082=3.3046grams"